Assume that you plan to take a housing loan with a tenor of 20 year. The loan has to be repaid in equal monthly installments. Considering that the loan amount is Rs. 50 lakhs and the interest rate on loan is 9% p.a., what would be the equated monthly installment (EMI)?


1
Expert's answer
2021-10-04T10:17:23-0400
"X=S(p+\frac{p}{(1-p)^N-1})"

"X=50(\frac{9}{12}+\frac{\frac{9}{12}}{(1+\frac{9}{12})^{480}-1})=3.75"